# How Do LLMs Use Their Depth?

### v1

@@ Abstract
footnotetext:    Correspondence to: Akshat Gupta <akshat.gupta@berkeley.edu>, Anna Ivanova <a.ivanova@gatech.edu>
Growing evidence suggests that large language models do not use their depth uniformly, yet we still lack a fine-grained understanding of their layer-wise prediction dynamics. In this paper, we trace the intermediate representations of several open-weight models during inference and reveal a structured and nuanced use of depth. Specifically, we propose a “Guess-then-Refine” framework that explains how LLMs internally structure their computations to make predictions. We first show that the top-ranked predictions in early LLM layers are composed primarily of high-frequency tokens, which act as statistical guesses proposed by the model early on due to the lack of appropriate contextual information. As contextual information develops deeper into the model, these initial guesses get refined into contextually appropriate tokens. Even high-frequency token predictions from early layers get refined $70\%$ of the time, indicating that correct token prediction is not “one-and-done”. We then go beyond frequency-based prediction to examine the dynamic usage of layer depth across three case studies. (i) Part-of-speech analysis shows that function words are, on average, the earliest to be predicted correctly. (ii) Fact recall task analysis shows that, in a multi-token answer, the first token requires more computational depth than the rest. (iii) Multiple-choice task analysis shows that the model identifies the format of the response within the first half of the layers, but finalizes its response only toward the end. Together, our results provide a detailed view of depth usage in LLMs, shedding light on the layer-by-layer computations that underlie successful predictions and providing insights for future works to improve computational efficiency in transformer-based models.
Code : https://github.com/akshat57/how-do-llms-use-their-depth
@@ 1 Introduction
The LogitLens framework (Nostalgebraist, 2020) provides some initial insights by using the final unembedding layer of a model to decode intermediate layer representations in the token space. This approach shows that, sometimes, generated tokens already begin to appear as top-ranked predictions at intermediate layers, sometimes long before the final layer. Another tool, TunedLens (Belrose et al., 2023), adapts the original LogitLens approach by learning a linear transformation between a model’s intermediate layer and the final layer, which enables tracking of token-level predictions without assuming that they lie in the same space. Here, we leverage TunedLens to detail and quantify token prediction patterns across LLM layers.
@@
Figure 1: Illustrative depictions of our Guess-then-Refine framework and complexity-aware depth use in LLMs. (a) Early layers elevate high-frequency tokens as statistical guesses, which deeper layers refine into context-appropriate predictions (Frequency-Conditioned Onset). (b) Multi-token fact recall: the first token of the answer requires the largest depth on average, while subsequent tokens appear at much shallower depths. (c) Option-constrained downstream tasks (e.g., True/False): early layers collect valid options into the top ranks, while later layers reason between top choices to generate final answer. Overall: LLMs act as early statistical guessers and late contextual integrators, exhibiting complexity-driven depth use.
@@
We propose a “Guess-then-Refine” framework for understanding model computations across layers, where models first create early guesses that get refined in subsequent layers. We first show that early-layer predictions in LLMs are largely composed of high-frequency tokens. For example, for Pythia6.9B, while the Top-10 most frequent tokens account for about 34% of the final predictions, the top-1 ranked early layers predictions contain the Top-10 tokens more than 75% times. We call this “Frequency-Conditioned Onset”, where LLMs tend to use corpus statistics while making candidate proposals in early layers. We then show that these early layer proposals are infact guesses, which get heavily refined in subsequent layers, where more than 80% of the early layer predictions get refined into into contextually appropriate generations by the final layer.
@@
We also uncover that LLMs use their depth dynamically based on the task at hand. (i) First, we show that during regular next token prediction, easier to predict tokens corresponding to punctuation and function words begin to get predicted correctly by early-to-mid LLM layers, whereas harder to predict tokens like nouns and verbs appear much later. (ii) We then look at multi-token fact recall task, where the model is asked factual questions whose answers span multiple tokens. We see that the second and third answer token predictions appear much earlier than the first, indicating a higher computational load associated with selecting the initial response direction. (iii) Finally, we analyze the layer-wise prediction dynamics when the models perform downstream tasks with a constrained set of option choices, such as answering multiple-choice (MCQ) or True/False questions. We show this task can be broken down into two steps: by the middle layers, the model arrives at the valid option choices as top-ranked predictions, and at the later layers, it deliberates between the valid options to reach the final answer. These experiments are indicative of Complexity-Aware Depth Use, where easier tasks require fewer layers of computation.
@@
Thus, through various experiments, we show that LLMs internally structure their computations following a “Guess-then-Refine” strategy, where they propose high-frequency tokens as top-ranked proposals in early layers, which later get modified into contextually appropriate tokens in deeper layers. We also show that LLMs are natural dynamic depth models and use their depth intelligently based on the complexity of the task. More complex predictions require larger depth, while easier computations finish earlier. We perform this analysis on four open-weight models - GPT2-XL (Radford et al., 2019), Pythia-6.9B (Biderman et al., 2023), Llama2-7B (Touvron et al., 2023) and Llama3-8B (Meta, 2024). This analysis is done using the TunedLens probe (Belrose et al., 2023) probe, which allows us to decode earlier layer representations with higher fidelity. We also perform ablations on TunedLens to confirm that our results represent information content in early layers.
@@ 2 Background : TunedLens
Intermediate activations in LLMs can be projected onto the vocabulary space using the unembedding matrix (Nostalgebraist, 2020). Let $h^{l}$ represent the hidden representation after layer $l$. LogitLens (Nostalgebraist, 2020) projects these representations in the token space according to the following equation:
@@
where $\texttt{Norm}_{f}$ represents the final normalization [^1] We use normalization to mean both LayerNorm Ba (2016) used in GPT-family of models and RMSNorm Zhang & Sennrich (2019) used in the Llama-family of models. applied before the application of the unembedding matrix $W_{U}\in\mathbb{R}^{d\times|V|}$; where $|V|$ represents the vocabulary size, and $d$ is the hidden dimension size of the model.
$\texttt{TunedLens}(h^{l})=\texttt{LogitLens}(A_{l}h^{l}+b_{l})$ (2)
The TunedLens probes are trained to minimize the KL-divergence between the final layer probability distribution and an intermediate layer probability distribution according to the following equation:
$\mathop{\mathrm{argmin}}_{A_{\ell},\,b_{\ell}}\mathbb{E}\left[D_{KL}\left(f_{\theta}(h^{l})\,\|\,\texttt{TunedLens}(h^{l})\right)\right]$ (3)
where $f_{\theta}(h)$ represents the probability distribution for input representation $h^{l}$ at the output of the model. Thus, TunedLens translators $(A_{\ell},\,b_{\ell})$ are trained to allow for better transferability between intermediate layers and the final unembedding matrix, which provide more faithful decoding of intermediate representations of the model, especially in the early layers (Belrose et al., 2023). This is why we opt to use TunedLens for for decoding the intermediate representations of LLMs.
3 Frequency-Conditioned Onset in Early LLM Layers
We begin by studying the top-ranked predictions across LLM layers during next token prediction using the TunedLens probe. We look at four open-weight models - GPT2-XL (Radford et al., 2019), Pythia-6.9B (Biderman et al., 2023), Llama2-7B (Touvron et al., 2023) and Llama3-8B (Meta, 2024). We use the TunedLens probes provided by the original authors for these models [^2] https://huggingface.co/spaces/AlignmentResearch/tuned-lens/tree/main.
Methodology: We divide vocabulary tokens of each model into four buckets by their frequency of occurrence in a corpus. For this, we take the tokenizer of each model and tokenize the English Wikipedia as distributed via the Hugging Face Datasets library (Lhoest et al., 2021), and record the frequency of each token. We use the following four buckets - (i) Top1-10 bucket, which contains the top 10 most frequent tokens in a corpus and accounts for approximately 23-25% of the corpus depending on the model, (ii) Top11-100 bucket, which contains the next 90 most frequent tokens and accounts for the next 16-20% of corpus tokens, (iii) Top101-1000 bucket, which contains the next 900 most frequent tokens and accounts for 16-21% of corpus tokens, and finally the (iv) Top1000+ bucket, which contains the remaining tokens and accounts for 35-40% of the corpus.
3.1 LLMs are Early Statistical Guessers
The top-ranked predictions at each intermediate layer classified into corresponding frequency buckets are shown in Figure 2. We discover that early layer predictions are heavily dominated by tokens belonging to the Top1-10 bucket. For example, more than 75% top-ranked tokens for Pythia-6.9B (Figure 2(a)) belong to the Top1-10 bucket at the first layer, while only 33% of the final predictions eventually made by the model lie in the Top1-10 bucket. Similarly for Llama3-8B (Figure 2(b)), the early layer predictions belong to the Top1-10 bucket for about 57% prefixes at the first layer, while the final predictions lie in the Top1-10 bucket only about 30% of times. A similar trend is observed in GPT2-XL and Llama2-7B, as shown in Appendix B.
@@ 3.1 LLMs are Early Statistical Guessers
This observation shows that LLMs follow a very specific pattern when making initial proposals, which can be explained by understanding the amount of information models have to work with in early layers when compared to deeper layers. In early layers, the model has incomplete contextual information about an input sentence, since the input has not been through enough attention layers to aggregate a complete contextual representation of the input. Additionally, in early layers, the model is also unable to access the factual knowledge stored in its parameters to make a correct predictions, since the knowledge stored in model parameters usually exists within the middle MLP layers of a model (Geva et al., 2020; Meng et al., 2022; Geva et al., 2023). In the absence of enough contextual information and learnt knowledge, the best strategy is to revert back to corpus statistics. To explain this, let us take the example of an extreme scenario where we do not have access to any information about the input prefix. In such a situation, if one had to guess the next token, picking a token from the Top1-10 bucket, which accounts for about 25% of the corpus, would be the best strategy as it maximizes the chances of being correct with a minimum set of tokens. Such an intuitive strategy naturally emerges out of the optimization pressures of the pre-training process.
@@
Thus, LLM make statistical guesses by promoting the most frequent tokens as top-ranked proposals in early layers. As contextual information develops deeper into the model, these guesses based on corpus statistics get refined into appropriate predictions, which is when lower frequency tokens begin to appear.
@@ 3.2 Massive Contextual Refinement in Later Layers
Indications for later layer refinements can already be seen in Figure 2, where the top-ranked tokens in early layers belong to the Top1000+ bucket for only about 10% of the prefixes, a number that increases by around 300% by the final layer. However, a clearer picture of the magnitude with which early layer proposals in LLMs get refined is shown in Figure 3. In this figure, the y-axis measures the percentage of top-ranked predictions at an intermediate layer that get modified by the final layer, bucketted into corresponding frequency buckets. For Pythia-6.9B (Figure 3(a)), while early layer decisions are dominated by the Top1-10 bucket tokens as shown in previous section, almost 80% of those predictions get modified by the final layer. This is also true for Llama3-8B (Figure 3(b)). Additionally, if a prediction at layer 1 lies in the Top1000+ bucket, it gets modified by the final layer with almost 100% probability. Results for the remaining models can be found in Appendix B. This shows that early layer proposals get modified in overwhelming numbers by the time they reach the final layer and indicates that refinement happens for both high- and low-frequency tokens.
@@
Thus, early LLM layers propose frequency-conditioned guesses that are highly non-permanent, where approximately 60–80% of early top-ranked guesses are eventually replaced. On the other hand, permanence rises with depth along with context integration. This means that onset of a token as top-ranked proposal, which is primarily based on corpus statistics, does not mean that the model has committed to a prediction, and future refinement based on information gathered in subsequent layers eventually decides final predictions.
(a) Pythia-6.9B
(b) Llama3-8B
Figure 3: Early-layer predictions are heavily revised. For each layer and frequency bucket, we show the fraction of top-1 ranked predictions that are overturned by the final layer (“decision flip rate”). We see that about 80% of layer-1 Top1–10 predictions change by the end, and this rises to 1̃00% for Top101–1000 and Top1000+ bucket tokens. As expected, flip rates drop with depth.
4 Complexity-Aware Depth Use in LLMs
In the previous section, we saw that LLMs propose initial guesses in early layers based on corpus statistics, which get refined in large numbers as more contextual information develops through the model. In this section, we go into more details of this guess-then-refine process across layers. Specifically, we show that LLMs are natural dynamic depth models that use their depth flexibly by performing easier to execute tasks (or subtasks) at shallower depths, while leaving more complicated tasks for later layers. We show this using three case studies: next-token prediction, fact recall, and fixed-response tasks. We perform all experiments for GPT2-XL, Pythia-6.9B, Llama2-7B and Llama3-8B. We present the results for Pythia-6.9B and Llama3-8B in the main paper and results for the remaining models can be found in Appendix C.
4.1 Case Study I: Depth Use by POS Category
Methodology: We first study the depth use strategy in LLMs during next-token prediction. Here, instead of looking at the top-ranked prediction at each layer, we track how the predicted token becomes top-ranked through the model depth. To do this, we send a prefix text as input to a model, determine the next predicted token, and then track the rank of the predicted token through the intermediate layers. The predicted token is classified into one of the following part-of-speech (POS) categories - determiners (DET), adpositions [^3] Adpositions include prepositions and postpositions (ADP), punctuations (PUNCT), adjectives (ADJ), verbs, and nouns. We plot the average layer at which a particular rank threshold is crossed by the predicted token at intermediate layers.
(a) Pythia-6.9B (POS)
(b) Llama3-8B (POS)
(c) Pythia-6.9B (Fact)
(d) Llama3-8B (Fact)
For example, if the input prefix sentence is “Today is a sunny” and the model predicts “day” as the generation, we first use Spacy [^4] https://spacy.io/ to classify the POS of the predicted word. Then, we use TunedLens to trace the predicted rank of “day” at different layers within the model. The first layer inside a model at which the predicted token crosses a particular rank threshold is recorded [^5] Rank progression in intermediate layers is not monotonic, and often predicted tokens can becomes top-ranked tokens at an intermediate layers and then increase in rank. However, here we record the first evidence of the predicted token crossing a particular rank threshold to track the first occurrence of the correct prediction in the “guess-then-refine” process..
We use POS to categorize generated tokens because they give us an indication of the role a generated word plays in a sentence. POS categories like DET and ADP contain function words that do not provide much meaning to a sentence, whereas other categories like adjectives, verbs, and nouns are made up of content words where the main information content of a sentence is present.
Results: Figures 4(a) and 4(b) show the average number of layers required for a predicted token to first cross a rank-threshold. Function words (DET and ADP) and punctuation (PUNCT) progress and reach rank-1 much faster than content words (ADJ, VERB, NOUN). For example, when the predicted word is a DET, it first reaches rank 1 on average around layer 5 for both Pythia-6.9B and Llama3-8B, while other content words categories reach rank-1 much deeper into a model, closer to layer 20.
This result is also consistent with the finding in Section 3: DET, ADP, and PUNCT contain primarily high-frequency tokens and are therefore predicted early in the model, at the initial guessing stage. Content words, however, must rely on contextual inference in middle layers to be guessed correctly.
Overall, Case Study I suggests that easier-to-predict tokens are guessed earlier in the model than harder-to-predict tokens. However, it can also be explained by the general token frequency effect, whereby frequent tokens happen to be predicted during the initial model guessing stage. To disentangle the ease-of-prediction effect from the token frequency effect, we turn to Case Study II.
4.2 Case Study II: Depth Use during Multi-Token Fact Recall
Methodology: In this case study, we track the predicted tokens through intermediate layers during fact recall using the MQuAKE dataset (Zhong et al., 2023), which contains a variety of single-token and multi-token facts. An example query that is input to the model in this dataset is - “The Statue of Liberty is located in” and the expected generated token is “New York City”. We only study internal computations in scenarios where the model predicts the correct answer. We also split the dataset into three different cases where the answer contains one, two or three tokens and look at how the rank of the predicted tokens proceed for each generated token of the answer. More information about the dataset statistics can be found in Appendix C.
Results: Figures 4(c) and 4(d) show the first layer at which the correct fact recall token crosses a given rank threshold. The first observation is that all computations for fact recall tokens happen much later than function word categories from Case Study I. While function tokens begin to appear around layer 5 in Pythia-6.9B and Llama3-8B, tokens corresponding to accurate fact recall on average begin to appear after layer 15. This means that recalling facts is not as easy for a model as generating DET or ADP categories and requires much larger depth. We also see that the first appearance of the single-token answer happens much earlier into a model (layer  15 for Pythia-6.9B, layer  20 for Llama3-8B) compared to the first token of the multi-token facts (layer  25 for both models). This shows that recalling multi-token facts, especially the first token of a multi-token fact, is harder for the model compared to a single token fact. Predicting a single token fact only requires accurate construction of a single token, while predicting a multi-token fact also requires a model to anticipate the future tokens that complete fact recall accurately, which may make the recalling multi-token facts harder.
Critically, while computing a multi-token fact, the earliest appearance of the first token in the multi-token fact requires the largest number of layers, but the subsequent tokens become top-ranked much sooner. For instance, for a 3-token fact, the first token on average emerges around layer 27 for Pythia-6.9B, whereas the subsequent second and third tokens begin to emerge around layers 20 and 12 respectively. This pattern represents another example of a graded use of model depth, where harder to do tasks, like predicting the first token in a multi-token fact, requires additional depth, while subsequent tokens that are conditionally easier to predict given the first token, appear at shallower depths. GPT2-XL and LLama3-8B follow similar trends, which can be seen in Appendix C.
@@ 4.2 Case Study II: Depth Use during Multi-Token Fact Recall
Figure 5: Mean rank of answer tokens during constrained-choice downstream tasks. As the model performs tasks like MMLU (response options: ABCD) and sentiment analysis ( response options: positive/negative), we track the mean rank of the options through the layers. Early layers of the model promote valid options choices as top-ranked candidates, whereas later layers reason between option choices (shown as zoomed-in figures in the top-right corner of each plot). This figur shows results for Pythia-6.9B. Results for other models and tasks follow similar patterns and can be found in Appendix C.1.
4.3 Case Study III: Depth Use during Downstream Tasks

### v2

@@ Abstract
Growing evidence suggests that large language models do not use their depth uniformly, yet we still lack a fine-grained understanding of their layer-wise prediction dynamics. In this paper, we trace the intermediate representations of several open-weight models during inference and reveal a structured and nuanced use of depth. Specifically, we propose a “Guess-then-Refine” framework that explains how LLMs internally structure their computations to make predictions. We first show that the top-ranked predictions in early LLM layers are composed primarily of high-frequency tokens, which act as statistical guesses proposed by the model due to the lack of contextual information. As contextual information develops deeper into the model, these initial guesses get refined into contextually appropriate tokens. We then examine the dynamic usage of layer depth through three case studies. (i) Multiple-choice task analysis shows that the model identifies appropriate options within the first half of the model and finalizes the response in the latter half. (ii) Fact recall task analysis shows that in a multi-token answer, the first token requires more computational depth than the rest. (iii) Part-of-speech analysis shows that function words are, on average, the earliest to be predicted correctly. To validate our results, we supplement probe-based analyses with causal manipulations in the form of activation patching and early-exiting experiments. Together, our results provide a detailed view of depth usage in LLMs, shedding light on the layer-by-layer computations that underlie successful predictions and providing insights for future works to improve computational efficiency in transformer-based models.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Illustrative depictions of our Guess-then-Refine framework and complexity-aware depth use in LLMs. (a) Early layers elevate high-frequency tokens as statistical guesses, which deeper layers refine into context-appropriate predictions (Frequency-Conditioned Onset). (b) Option-constrained downstream tasks (e.g., True/False): early layers collect valid options into the top ranks, while later layers reason between top choices to generate final answer. (c) Multi-token fact recall: the first token of the answer requires the largest depth on average, while subsequent tokens appear at much shallower depths. Overall: LLMs act as early statistical guessers and late contextual integrators, exhibiting complexity-driven depth use.
@@
We present a “Guess-then-Refine” framework to answer these questions and understand model computations across layers. Under this framework, we show that models first create early guesses that get refined in subsequent layers. We first show that early-layer predictions in LLMs are largely composed of high-frequency tokens, like determiners or prepositions. For example, for Pythia6.9B, while the Top-10 most frequent tokens account for about 34% of the final predictions, the top-1 ranked early layer predictions contain the Top-10 tokens more than 75% times. We call this “Frequency-Conditioned Onset”, where LLMs tend to use corpus statistics while making candidate proposals in early layers. We then show that these early layer proposals are guesses which get heavily refined in subsequent layers, where more than 80% of the early layer predictions get refined into into contextually appropriate generations by the final layer. We leverage TunedLens (Belrose et al., 2023) to quantify token prediction patterns across LLM layers.
@@
We also show that LLMs use their depth dynamically based on the task at hand through three case studies. (i) First, we analyze the layer-wise prediction dynamics when the models perform downstream tasks with a constrained set of option choices, such as answering multiple-choice (MCQ) or True/False questions. We uncover that the model performs this task in two steps with a clear phase transition: it first collects the valid option choices as top-ranked predictions by the middle layers, and in later layers the model deliberates between the valid options to reach the final answer. We also accompany this case study with activation patching experiment showing causal evidence of a clear phase transition happening as the model performs such downstream tasks. (ii) We then look at multi-token fact recall task, where the model is asked factual questions whose answers span multiple tokens. We see that the second and third answer token predictions appear much earlier than the first, indicating a higher computational load associated with selecting the initial response direction. (iii) Finally, we show that during regular next token prediction, easier to predict tokens corresponding to function words and punctuation begin to get predicted correctly by early-to-mid LLM layers, whereas harder to predict tokens like nouns and verbs appear much later. The above two case studies are also accompanied by early exiting experiments. Together, these experiments are indicative of Complexity-Driven Depth Use, where easier tasks require fewer layers of computation.
@@
We perform this analysis on four open-weight models - GPT2-XL (Radford et al., 2019), Pythia-6.9B (Biderman et al., 2023), Llama2-7B (Touvron et al., 2023) and Llama3-8B (Meta, 2024). The experiments are done using the TunedLens probe probe, which allows us to decode earlier layer representations with higher fidelity. We also provide independent evidence with causal ablations and early exiting experiments to confirm our findings. Addionally, we also perform ablations on TunedLens to confirm that our results represent information content in early layers.
@@ 2 Background : TunedLens
Intermediate activations in LLMs can be projected onto the vocabulary space using the unembedding matrix (Nostalgebraist, 2020). Let $h^{l}$ represent the hidden representation after layer $l$. LogitLens (Nostalgebraist, 2020) projects these representations in the token space as follows:
@@
where $\texttt{Norm}_{f}$ represents the final normalization applied before the application of the unembedding matrix $W_{U}\in\mathbb{R}^{d\times|V|}$; where $|V|$ represents the vocabulary size, and $d$ is the hidden dimension size of the model.
$\texttt{TunedLens}(h^{l})=\texttt{LogitLens}(A_{l}h^{l}+b_{l})$ (2)
The TunedLens probes are trained to minimize the KL-divergence between the final layer and an intermediate layer probability distribution. Thus, TunedLens translators $(A_{\ell},\,b_{\ell})$ are trained to allow for better transferability between intermediate layers and the final unembedding matrix, which provide more faithful decoding of intermediate representations of the model, especially in the early layers (Belrose et al., 2023).
3 Frequency-Conditioned Onset in Early LLM Layers
We begin by studying the top-ranked predictions across LLM layers during next token prediction using TunedLens. We look at four open-weight models - GPT2-XL (Radford et al., 2019), Pythia-6.9B (Biderman et al., 2023), Llama2-7B (Touvron et al., 2023) and Llama3-8B (Meta, 2024). We use the TunedLens probes provided by the original authors for these models (AlignmentResearch, 2026).
Methodology: We divide vocabulary tokens of each model into four buckets by their frequency of occurrence. For this, we take the tokenizer of each model and tokenize the English Wikipedia dataset (Lhoest et al., 2021), and record the frequency of each token. We use the following four buckets - (i) Top1-10 bucket, which contains the top 10 most frequent tokens in a corpus and accounts for approximately 23-25% of the corpus depending on the model, (ii) Top11-100 bucket, which contains the next 90 most frequent tokens and accounts for the next 16-20% of corpus tokens, (iii) Top101-1000 bucket, which contains the next 900 most frequent tokens and accounts for 16-21% of corpus tokens, and finally the (iv) Top1000+ bucket, which contains the remaining tokens and accounts for 35-40% of the corpus.
(a) Pythia-6.9B
(b) Llama3-8B
Figure 3: Early-layer predictions are heavily revised. For each layer and frequency bucket, we show the fraction of top-1 ranked predictions that are overturned by the final layer (“decision flip rate”). We see that about 80% of layer-1 Top1–10 predictions change by the end, and this rises to about 100% for Top101–1000 and Top1000+ bucket tokens. As expected, flip rates drop with depth.
3.1 LLMs are Early Statistical Guessers
The top-ranked predictions at each intermediate layer classified into corresponding frequency buckets are shown in Figure 2. We discover that early layer predictions are heavily dominated by tokens belonging to the Top1-10 bucket. For example, more than 75% top-ranked tokens for Pythia-6.9B (Figure 2(a)) belong to the Top1-10 bucket at the first layer, while only 33% of the final predictions eventually made by the model lie in the Top1-10 bucket. A similar trend can be seen for Llama3-8B (Figure 2(b)) and GPT2-XL and Llama2-7B, as shown in Appendix B.
@@ 3.1 LLMs are Early Statistical Guessers
This behavior can be explained by understanding the amount of information available in early layers when compared to deeper layers. In early layers, the model has incomplete contextual information about the input sentence, since the input has not been through enough attention layers to aggregate this information. Additionally, in early layers, the model is also unable to access the factual knowledge stored in its parameters, since the stored knowledge lies in the middle MLP layers (Geva et al., 2020; Meng et al., 2022; Geva et al., 2023). In the absence of enough contextual information and learnt knowledge, the best strategy is to revert back to corpus statistics. To explain this, let us take the example of an extreme scenario where we do not have access to any information about the input prefix. In such a situation, if one had to guess the next token, picking a token from the Top1-10 bucket, which accounts for about 25% of the corpus, would be the best strategy as it maximizes the chances of being correct with a minimum set of tokens. Such an intuitive strategy naturally emerges out of the optimization pressures of the pre-training process!
@@
Thus, LLMs make statistical guesses by promoting the most frequent tokens as top-ranked proposals in early layers. As contextual information develops deeper into the model, these guesses based on corpus statistics get refined into appropriate predictions, which is when lower frequency tokens begin to appear.
@@ 3.2 Massive Contextual Refinement in Later Layers
Indications for later layer refinements can already be seen in Figure 2, where the top-ranked tokens in early layers belong to the Top1000+ bucket for only about 10% of the prefixes, a number that increases by around 300% by the final layer. However, here we consider a more specific metric for refinement. We define refinement with the “decision flip rate”, which measures the percentage of top-1 ranked tokens at an intermediate layer $l$ that are not the same as the final prediction.
@@
Figure 3 shows a clearer picture of the magnitude with which early layer proposals in LLMs get refined to final predictions. For Pythia-6.9B (Figure 3(a)), while early layer decisions are dominated by the Top1-10 bucket tokens as shown in previous section, almost 80% of those predictions get modified by the final layer. This is also true for Llama3-8B (Figure 3(b)). Additionally, if a prediction at layer-1 lies in the Top1000+ bucket, it gets modified by the final layer with almost 100% probability. Results for the remaining models can be found in Appendix B. This shows that early layer proposals get modified in overwhelming numbers by the time they reach the final layer and indicates that refinement happens for both high- and low-frequency tokens.
@@
Thus, early LLM layers propose frequency-conditioned guesses that are highly non-permanent, where approximately 60–80% of early top-ranked guesses are eventually replaced. On the other hand, permanence rises with depth along with context integration.
@@
Figure 4: Mean rank of answer tokens during constrained-choice downstream tasks. As the model performs tasks like MMLU (response options: ABCD) and sentiment analysis ( response options: positive/negative), we track the mean rank of the options through the layers. Early layers of the model promote valid options choices as top-ranked candidates, whereas later layers reason between option choices (shown as zoomed-in figures in the top-right corner of each plot). This figure shows results for Pythia-6.9B. Results for other models and tasks follow similar patterns and can be found in Appendix C.1. (c) Activation patching experiments, where we take activations from an SST forward pass and patch it during the forward pass of MMLU. This figure shows in early layers, the model is still able to find one of MMLU options (A/B/C/D), whereas in later layers the generated answer is an SST option (positive negative). This shows causal evidence of the two phase behavior shown in subfigure (a) and (b). Additionaly, the phase-transition layers are predictions made by TunedLens, which very cleanly match actual model behavior.
4 Complexity-Driven Depth Use in LLMs
In the previous section, we saw that LLMs propose initial guesses in early layers based on corpus statistics, which get refined in large numbers as more contextual information develops through the model. In this section, we go into more details of this guess-then-refine process across layers. Specifically, we show that model depth in LLMs gets used flexibly, where easier to execute tasks (or subtasks) finish at shallower depths, while more complicated tasks are left for later layers. We show this using three case studies: fixed-option question answering, fact recall and next-token prediction. We also perform causal ablations and early-exiting experiments to validate our findings.
4.1 Case Study I: Depth Use during Downstream Tasks